THE MILLIMETER/SUBMILLIMETER SPECTRUM OF BaSH(X̃^{2}A′): EVEN FURTHER INVESTIGA-TIONS OF THE METAL-SULFUR BOND

\underline{A. JANCZYK} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of BaSH (X̃^{2}A′) and one of its isotopomers, BaSD, were recorded using millimeter/submillimeter wave direct absorption techniques in the range 320-364 GHz. This radical was produced by reacting barium vapor with H_{2}S in the presence of a d.c. discharge. Ten transitions, ranging from J=66 → 67 to 75 → 76, were measured for BaSH, as well as 5 transitions for the deuterium isotopomer. The spectra exhibit complex K_{a} ladder structure for both molecules, consistent with a bent geometry. Spin- rotation splittings of approximately 54 MHz were also observed in every transition. Rotational and fine structure constants have been determined for both isotopomers, as well as an r_{0} structure for BaSH. This study shows that barium hydrosulfide is bent with an angle of Θ = 88^{∘}, and therefore is significantly different from linear BaOH. This structural change indicates more covalent bonding character in the BaSH, relative to its hydroxide analog.